OpenStudy (anonymous):

STUDENTS AT A UNIVERSITY CAMPUS 22 or younger =44.7 23–25 years =25.2 26–35 years= 20.1 36 or older = 10.0 The table above shows the percent of students in different age groups at a university campus. The total number of students at the campus is 8,987. Approximately how many of the students at the campus are of age 25 years or younger?

OpenStudy (anonymous):

so it would be 69.9 percent of 8.987 but how do i figure this out????

OpenStudy (anonymous):

The total number of students is 100% of the total. The amount of students aged 22 or younger is 44.7% of the total. The amount of students aged 23-25 years is 25.2% of the total. Therefore, the percentage of students aged 25 or less is the sum of those aged 23-25 and those under 23. Add them together.

OpenStudy (anonymous):

ray got part of the answer. however he only gave you a percentage and you need a number of students.

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

if you still need the final number of students with an age less than 25. add up the percentages for ages less than or equal to 25. 22 or younger =44.7 23–25 years =25.2 25.2 + 44.7 = 69.9 % multiply this by the total number of students to find the number of students whose age is less than or equal to 25 years of age.

OpenStudy (anonymous):

by the way, is this for math stats or psych/business stats.

OpenStudy (anonymous):

for math stats @tomo

OpenStudy (anonymous):

i am still unsure about what to do :o

Directrix (directrix):

@helpmexoxo >>so it would be 69.9 percent of 8.987 If you are asking how to take 69.9% of 8,987, then recall that 69.9% = 69.9/100 = .699. Then, multiply .699 times 8,987. That product should give the number of students age 25 or younger.